# A Survey on LLM-as-a-Judge

### v2

@@ 1. Introduction
Recently, Large Language Models (LLMs) have achieved remarkable success across numerous domains, ranging from technical fields to the humanities and social sciences. Building on their success, the concept of using LLMs as evaluators—commonly referred to as ”LLM-as-a-Judge” (Zheng et al., 2023b)—has gained significant attention, where LLMs are tasked with determining whether something falls within the scope of a given rule. This growing interest stems from LLMs’ ability to mimic human-like reasoning and thinking processes, enabling them to take on roles traditionally reserved for human experts while offering a cost-effective solution that can be effortlessly scaled to meet increasing evaluation demands. For instance, employing LLM-as-a-Judge in the academic peer-review [^1] https://blog.iclr.cc/2024/10/09/iclr2025-assisting-reviewers/ process can help handle the rapid increase in submissions while maintaining expert-level judgment.
Before the era of LLMs, achieving a balance between comprehensive and scalable evaluation had posed a persistent challenge. On the one hand, commonly used subjective methods like expert-driven assessments integrate holistic reasoning and nuanced contextual understanding, making them the gold standard in comprehensiveness  (Shi et al., 2024a; Gao et al., 2023). However, these approaches are expensive, challenging to scale, and susceptible to inconsistency. On the other hand, objective assessment methods like automatic metrics offer strong scalability and consistency. For example, tools such as BLEU (Papineni et al., 2002) or ROUGE (Lin, 2004) can quickly score machine-generated translations or summaries against reference texts without human intervention. Yet these metrics, which heavily rely on surface-level lexical overlaps, often fail to capture deeper nuances of meaning, coherence, or logical reasoning. As a result, they perform poorly when evaluating tasks like open-ended story generation or complex instructional texts (Schluter, 2017). With the rise of advanced LLM techniques, the “LLM-as-a-Judge” has emerged as a potential idea to combine the strengths of above two ways of evaluation. LLMs have been proven to merge the scalability of automatic methods with the nuanced, context-sensitive reasoning found in expert judgments (Zheng et al., 2023b; Wang et al., 2023d; Zhu et al., 2023a; Li et al., 2023b; Chen et al., 2024d). Moreover, LLMs might become flexible enough to handle multimodal inputs citemllm-as-a-judge under decent prompt learning or fine-tuning (Khattak et al., 2023), which is a unique advantage compared with traditional methods. These advantages suggest that the LLM-as-a-Judge approach could serve as a novel and broadly applicable paradigm for addressing complex and open-ended evaluation problems.
@@
To address these challenges, this paper provides a systematic review of research on LLM-as-a-Judge, offering a comprehensive overview of the field while exploring strategies for building reliable LLM-as-a-Judge systems. We begin by defining LLM-as-a-Judge through both formal and informal definitions, answering the foundational query: ’What is LLM-as-a-Judge?’ Next, we categorize existing methods and approaches for its use, exploring how LLM-as-a-Judge can be effectively implemented in practice. Following this, we tackle the critical question: How to build reliable LLM-as-a-Judge systems? To answer this, we explore two core aspects: (1) strategies to enhance the reliability of LLM-as-a-Judge systems and (2) methodologies for evaluating the reliability of these systems. For the first aspect, we review key strategies to optimize the performance of LLM-as-a-Judge for diverse evaluation tasks. For the second aspect, we examine the metrics, datasets, and methodologies used to evaluate the performance of LLM-as-a-Judge systems, highlighting potential sources of bias and methods for their mitigation. Building on this foundation, we propose a novel benchmark specifically designed for evaluating LLM-as-a-Judge systems. Additionally, we explore practical application scenarios, identify challenges unique to each context, and propose solutions to address these issues. Finally, we discuss future research directions, emphasizing key areas for improving the reliability, scalability, and applicability of LLM-as-a-Judge systems.
The rest of this survey is organized as Figure 1. Section 2 provides an overview of the LLM-as-a-Judge field, including its definitions and categorization of existing methods.For a quick guide on implementing an LLM-as-a-Judge for specific scenarios, you can find answers in Quick Practice (2.5). Strategies for enhancing and evaluating the reliability of LLM-as-a-Judge systems are discussed in Sections 3 and 4, respectively. Section 6 explores practical applications, while Sections 7 and 8 address challenges and outline future research directions. Finally, Section 9 presents our conclusions.
@@ 2. Background and Method
The ability of LLMs to mimic human reasoning and evaluate specific inputs against a set of predefined rules has paved the way for ”LLM-as-a-Judge.” Their scalability, adaptability, and cost-effectiveness make them well-suited for managing a growing number of evaluative tasks that were traditionally done by humans. These abilities are key in utilizing LLMs flexibly across various evaluation scenarios and objectives. The adoption of LLM-as-a-Judge has progressed rapidly. Initially, the primary focus of LLMs was on language generation and comprehension. With advancements in training paradigms like Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022a), LLMs became more aligned with human values and reasoning processes. This alignment has allowed LLMs to transition from generative tasks to evaluative roles. At its core, LLM-as-a-Judge refers to the use of LLMs to evaluate objects, actions, or decisions based on predefined rules, criteria, or preferences. It encompasses a broad spectrum of roles, including: Graders (Trung et al., 2024), Evaluators/Assessors (Li et al., 2024b; Zhang et al., 2024b), Critics (Ke et al., 2024; Xiong et al., 2024; Putta et al., 2024), Verifiers (Ling et al., 2024; Shinn et al., 2023; Wang et al., 2024c), Examiners (Bai et al., 2023a), Reward/Ranking Models (Shi et al., 2024b; Sun et al., 2023a; Luo et al., 2023a; Yuan et al., 2023), etc.
@@
  • $\mathcal{E}$: The final evaluation obtained from the whole LLM-as-a-Judge process in the expected manner. It could be a score, a choice, or a sentence, etc.
@@
The formulation of LLM-as-a-Judge reflects that LLM is a type of auto-regressive generative model, which generates subsequent content based on the context then obtain target evaluation from it. It illustrates how we utilize LLM for evaluation tasks, encompassing input design, model selection and training, as well as output post-processing. The basic approaches of implementing LLM-as-a-Judge can be classified according to the formulation: In-Context Learning, Model Selection, Post-processing Method and Evaluation Pipeline, which concluded in Figure 2. By following this pipeline, one can build a basic LLM-as-a-Judge for evaluation. A quick practice guide is available in section 2.5.
@@ 2.1.1. Generating scores
Figure 3. The template for Likert scale scoring from Gao et al. (2023).
@@ 2.1.3. Conducting pairwise comparisons
Figure 5. The template for pairwise comparison from Gao et al. (2023)
@@ 2.4. Evaluation Pipeline
After completing the three processes, we obtain the final evaluation. From input to output, these steps collectively form the LLM-as-a-Judge evaluation pipeline, as illustrated in Figure 2. This pipeline is commonly applied in three scenarios: LLM-as-a-Judge for LLMs, LLM-as-a-Judge for data, and LLM-as-a-Judge for agents.
2.4.1. LLM-as-a-Judge for model
It is universally known that the best way to evaluate LLMs is human judgment, but collecting human annotations can be costly, time-consuming, and laborious (Ouyang et al., 2022b; Zheng et al., 2023d). Using strong LLMs (usually closed-source ones, e.g., GPT-4, Claude, ChatGPT) as an automated proxy for assessing LLMs has become a natural choice (Zhou et al., 2023). With appropriate prompt design, the quality of evaluation and agreement to human judgment can be promising (Dubois et al., 2023; Zheng et al., 2023d; Zhang et al., 2023d; Wang et al., 2023b). However, the cost concern still exists when calling the APIs of these proprietary models, especially when there is a frequent need for model validation on large-scale data. Moreover, closed-source LLM-as-a-Judge leads to low reproducibility due to potential changes in models behind the API. Some recent works have started to make attempts for open-source alternatives. SelFee (Ye et al., 2023) collects generations, feedback, and revised generations from ChatGPT and fine-tunes LLaMA models to build a critique model. Shepherd (Wang et al., 2023c) trains a model that can output critiques for single-response with the data of feedback from online communities and human annotation. PandaLM (Wang et al., 2023d) trains a model to conduct pairwise comparison for LLM Instruction Tuning Optimization, and Zheng et al. (2023d) also fine-tune Vicuna (Touvron et al., 2023b) on a 20K pairwise comparison dataset to explore the potential of open-source models as a more cost-friendly proxy.
Recent advancements in using Large Multimodal Models (LMMs) as evaluators have showcased their potential to perform complex judgment tasks in vision-language scenarios. Proprietary models like GPT-4V and GPT-4o have been pivotal in benchmarks such as detailed captioning and visual chats, utilizing both pointwise and pairwise evaluation methods (Liu et al., 2023b; Zhang et al., 2023b; Lu et al., 2024). Open-source alternatives have emerged, with Prometheus-Vision (Lee et al., 2024a) being the first vision-language model specifically trained to act as an evaluator for user-designed scoring criteria. While Prometheus-Vision introduced the concept of open-source evaluators with a focus on specialized tasks, it remains limited to predefined criteria. In contrast, LLaVA-Critic (Xiong et al., 2024), another open-source innovation, expands the scope by serving as a generalist evaluator. Trained on diverse and detailed datasets, LLaVA-Critic provides robust scoring and preference learning, closely aligning with human and proprietary evaluations. These models mark significant progress in democratizing and enhancing multimodal evaluation tools.
@@ 2.4.2. LLM-as-a-Judge for data
Data annotation generally refers to the labeling or generating of raw data with relevant information, which could be used for improving the efficacy of machine learning models. The process, however, is labor-intensive and costly. The emergence of LLMs presents an unprecedented opportunity to automate the complicated process of data annotation by LLM-as-a-Judge. Most of the data need to be evaluated by LLM-as-a-Judge is generated by models, or large-scale crawled data. Language models first conduct supervised fine-tuning to imitate how to align with human instructions (Wang et al., 2023a; Taori et al., 2023). After that, reinforcement learning techniques have been explored to align language models with human preferences (Ouyang et al., 2022c; Ramamurthy et al., 2023). The most successful way is applying a RLHF framework (Ouyang et al., 2022c) via training a reward model on human feedback and using PPO (Schulman et al., 2017) to obtain the policy model for language generation. However, in practices, the PPO training paradigm is complex in coding and hyper-parameter tuning while it needs four models that are hard for training. This motivates us to explore simpler and more straightforward methods to align language models with human preferences. This involves how to use LLM-as-a-Judge to evaluate whether different responses are aligned with human preferences. For example,  (Yuan et al., 2023; Dong et al., 2023) use general LLM (ChatGPT) to get better alignment with human preferences. The Aplaca prompts (Taori et al., 2023) is used as sampling queries to different models generate responses. And these data was evaluated by LLM-as-a-Judge to obtain human preference scores (reward score) to train a new language model. Other works would like to use Supervised Fine-Tuning (SFT) model itself as evaluator, like generating better-aligned datasets for SFT including hindsight-modified prompts (Zhang et al., 2023a; Liu et al., 2023d) and principle-driven self-alignment (Sun et al., 2023c).
@@ 3. Improvement Strategy
When directly utilizing LLMs to conduct evaluation tasks such as scoring, selection, pairwise comparison or ranking, the inherent biases of LLMs like length bias, positional bias and concreteness bias(Park et al., 2024) will lead to poor evaluation results. Addressing these inherent biases and improving the overall evaluation performance of LLMs is a critical challenge for applying LLMs as evaluators. In this section, we introduce three improvement strategy aimed at enhancing the evaluation performance of LLM-as-a-judge: design strategy of evaluation prompts (in-context learning based), improvement strategy of LLMs’ evaluation capabilities (model based), and optimization strategy of final evaluation results (post-processing based). Our categorization is based on the formal definition of LLM-as-Evaluator in Section 2, focusing on enhancing the evaluation effectiveness by targeting three key phases of the process: the context $\mathcal{C}$, the abilities of LLMs themselves $\mathcal{P_{LLM}}$ and the post processing $\leftarrow$ to obtain the final results $\mathcal{E}$
@@
Output Forms, text width=6em [Constraining outputs in structured formats: G-Eval (Liu et al., 2023a) / DHP (Wang et al., 2024d) / LLM-EVAL (Lin and Chen, 2023)] [Providing evaluations with explanations: CLAIR (Chan et al., 2023) / FLEUR (Lee et al., 2024b)] ] ] [Improvement Strategy
@@
Based on Feedbacks, text width=7em [INSTRUCTSCORE (Xu et al., 2023) / JADE (Zhang et al., 2023c)] ] ] [Optimization Strategy
@@
Evaluation Results, text width=7.5em [Summarize by multiple rounds: Sottana et al. (Sottana et al., 2023) / PsychoBench (Huang et al., 2023) / Auto-J (Li et al., 2023b)] [Vote by multiple LLMs: CPAD (Liu et al., 2023e) / Bai et al. (Bai et al., 2023b)] ] [Direct Optimization of
LLMs’ Outputs, text width=7.5em [Score smoothing: FLEUR (Lee et al., 2024b) / G-Eval (Liu et al., 2023a) / DHP (Wang et al., 2024d)] [Self validation: TrueTeacher (Gekhman et al., 2023)] ] ] ]

### v3

@@ 1. Introduction
Recently, Large Language Models (LLMs) have achieved remarkable success across numerous domains, ranging from technical fields to the humanities and social sciences. Building on their success, the concept of using LLMs as evaluators—commonly referred to as ”LLM-as-a-Judge” (Zheng et al., 2023b)—has gained significant attention, where LLMs are tasked with determining whether something falls within the scope of a given rule (Kant, 1790, 1781). This growing interest stems from LLMs’ ability to mimic human-like reasoning and thinking processes, enabling them to take on roles traditionally reserved for human experts while offering a cost-effective solution that can be effortlessly scaled to meet increasing evaluation demands. For instance, employing LLM-as-a-Judge in the academic peer-review [^1] https://blog.iclr.cc/2024/10/09/iclr2025-assisting-reviewers/ process can help handle the rapid increase in submissions while maintaining expert-level judgment.
Before the era of LLMs, achieving a balance between comprehensive and scalable evaluation had posed a persistent challenge. On the one hand, commonly used subjective methods like expert-driven assessments (Shi et al., 2024; Gao et al., 2023a) integrate holistic reasoning and nuanced contextual understanding, making them the gold standard in comprehensiveness. However, these approaches are expensive, challenging to scale, and susceptible to inconsistency. On the other hand, objective assessment methods like automatic metrics offer strong scalability and consistency. For example, tools such as BLEU (Papineni et al., 2002) or ROUGE (Lin, 2004) can quickly score machine-generated translations or summaries against reference texts without human intervention. Yet these metrics, which heavily rely on surface-level lexical overlaps, often fail to capture deeper nuances, leading to poor performance in tasks like story generation or instructional texts (Schluter, 2017). As a solution to this persistent dilemma, “LLM-as-a-Judge” has emerged as a promising idea to combine the strengths of the above two evaluation methods, merging the scalability of automatic methods with the nuanced, context-sensitive reasoning found in expert judgments (Zheng et al., 2023b; Wang et al., 2023d; Zhu et al., 2023a; Li et al., 2023b; Chen et al., 2024e). Moreover, LLMs might become flexible enough to handle multimodal inputs (Chen et al., 2024d) under decent prompt learning or fine-tuning (Khattak et al., 2023). These advantages suggest that the LLM-as-a-Judge approach could serve as a novel and broadly applicable paradigm for addressing complex and open-ended evaluation problems.
@@
To address these challenges, this paper provides a systematic review of research on LLM-as-a-Judge, offering a comprehensive overview of the field while exploring strategies for building reliable LLM-as-a-Judge systems. We begin by defining LLM-as-a-Judge through both formal and informal definitions, answering the foundational question: ”What is LLM-as-a-Judge?” Next, we categorize existing methods and approaches, exploring ”How to use LLM-as-a-Judge?”. Following this, to tackle the critical question: ”How to build reliable LLM-as-a-Judge systems?”, we explore two core aspects: (1) strategies to enhance the reliability of LLM-as-a-Judge systems and (2) methodologies for evaluating the reliability of these systems. For the first aspect, we review key strategies to optimize the performance of LLM-as-a-Judge. For the second aspect, we examine the metrics, datasets, and methodologies used to evaluate LLM-as-a-Judge systems, highlighting potential sources of bias and methods for their mitigation. Building on this, we introduce a novel benchmark specifically designed for evaluating LLM-as-a-Judge systems. Additionally, we explore practical application scenarios and identify challenges unique to each context. Finally, we discuss future research directions, emphasizing key areas for improving reliability, scalability, and applicability.
The rest of this survey is organized as Figure 1. Section 2 provides an overview of the LLM-as-a-Judge field, including its definitions and categorization of existing methods. For a quick guide on the implementation of an LLM as a judge for specific scenarios, you can find answers in Quick Practice (2.5). Strategies for enhancing and evaluating the reliability of LLM-as-a-Judge systems are discussed in Sections 3,  4, and  5, respectively. Section 6 explores practical applications, while Sections 7 and 8 address challenges and outline future research directions. Finally, Section 9 presents our conclusions.
@@ 2. Background and Method
The ability of LLMs to mimic human reasoning and evaluate specific inputs against a set of predefined rules has paved the way for ”LLM-as-a-Judge.” Their scalability, adaptability, and cost-effectiveness make them well-suited for managing a growing number of evaluative tasks that were traditionally done by humans. These abilities are key in utilizing LLMs flexibly across various evaluation scenarios and objectives. The adoption of LLM-as-a-Judge has progressed rapidly. Initially, the primary focus of LLMs was on language generation and comprehension. With advancements in training paradigms like Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022a), LLMs became more aligned with human values and reasoning processes. This alignment has allowed LLMs to transition from generative tasks to evaluative roles. At its core, LLM-as-a-Judge refers to the use of LLMs to evaluate objects, actions, or decisions based on predefined rules, criteria, or preferences. It encompasses a broad spectrum of roles, including: Graders (Trung et al., 2024; Dong et al., 2023), Evaluators/Assessors (Li et al., 2024b; Zhang et al., 2024b), Critics (Ke et al., 2024; Xiong et al., 2024; Putta et al., 2024), Verifiers (Ling et al., 2024; Shinn et al., 2023; Wang et al., 2024c), Examiners (Bai et al., 2023a), Reward/Ranking Models (Yang et al., 2024b; Sun et al., 2023a; Luo et al., 2023a; Yuan et al., 2023), etc.
@@
  • $\mathcal{E}$: The final evaluation obtained from the whole LLM-as-a-Judge process in the expected manner. It could be a score, a choice, a label or a sentence, etc.
@@
The formulation of LLM-as-a-Judge reflects that LLM is a type of auto-regressive generative model, which generates subsequent content based on the context and then obtains target evaluation from it. It illustrates how we utilize LLM for evaluation tasks, encompassing input design, model selection, and training, as well as output post-processing. The basic approaches of implementing LLM-as-a-Judge can be classified according to the formulation: In-Context Learning, Model Selection, Post-processing Method, and Evaluation Pipeline, which concluded in Figure 2. By following this pipeline, one can build a basic LLM-as-a-Judge for evaluation. A quick practice guide is available in section 2.5.
@@ 2.1.1. Generating scores
Figure 3. The template for Likert scale scoring from Gao et al. (2023a).
@@ 2.1.3. Conducting pairwise comparisons
Figure 5. The template for pairwise comparison from Gao et al. (2023a)
@@ 2.4. Evaluation Pipeline
After completing the three processes, we obtain the final evaluation $\mathcal{E}$. From input to output, these steps collectively form the LLM-as-a-Judge evaluation pipeline, as illustrated in Figure 2. This pipeline is commonly applied in four scenarios showed in Figure 7: LLM-as-a-Judge for models, LLM-as-a-Judge for data, LLM-as-a-Judge for agents, and LLM-as-a-Judge for reasoning/thinking.
Figure 7. Four typical scenarios using LLM-as-a-Judge evaluation pipeline.
2.4.1. LLM-as-a-Judge for Models
It is universally known that the best way to evaluate LLMs is human judgment, but collecting human annotations can be costly, time-consuming, and laborious (Ouyang et al., 2022b; Zheng et al., 2023d). Using strong LLMs (usually closed-source ones, e.g., GPT-4, Claude, ChatGPT) as an automated proxy for assessing LLMs has become a natural choice (Zhou et al., 2023). With appropriate prompt design, the quality of evaluation and agreement to human judgment can be promising (Dubois et al., 2023; Zheng et al., 2023d; Zhang et al., 2023c; Wang et al., 2023b). However, the cost concern still exists when calling the APIs of these proprietary models, especially when there is a frequent need for model validation on large-scale data. Moreover, closed-source LLM-as-a-Judge leads to low reproducibility due to potential changes in models behind the API. Some recent works have started to make attempts for open-source alternatives. SelFee (Ye et al., 2023) collects generations, feedback, and revised generations from ChatGPT and fine-tunes LLaMA models to build a critique model. Shepherd (Wang et al., 2023c) trains a model that can output critiques for single-response with the data of feedback from online communities and human annotation. PandaLM (Wang et al., 2023d) trains a model to conduct pairwise comparison for LLM Instruction Tuning Optimization, and Zheng et al. (2023d) also fine-tune Vicuna (Touvron et al., 2023b) on a 20K pairwise comparison dataset to explore the potential of open-source models as a more cost-friendly proxy.
@@ 2.4.2. LLM-as-a-Judge for Data
Data annotation generally refers to the labeling or generating of raw data with relevant information, which could be used for improving the efficacy of machine learning models. The process, however, is labor-intensive and costly. The emergence of LLMs presents an unprecedented opportunity to automate the complicated process of data annotation by LLM-as-a-Judge. Most of the data need to be evaluated by LLM-as-a-Judge is generated by models, or large-scale crawled data. Language models first conduct supervised fine-tuning to imitate how to align with human instructions (Wang et al., 2023a; Taori et al., 2023). After that, reinforcement learning techniques have been explored to align language models with human preferences (Ouyang et al., 2022c; Ramamurthy et al., 2023). The most successful way is applying a RLHF framework (Ouyang et al., 2022c) via training a reward model on human feedback and using PPO (Schulman et al., 2017) to obtain the policy model for language generation. However, in practices, the PPO training paradigm is complex in coding and hyper-parameter tuning while it needs four models that are hard for training. This motivates us to explore simpler and more straightforward methods to align language models with human preferences. This involves how to use LLM-as-a-Judge to evaluate whether different responses are aligned with human preferences. For example,  (Yuan et al., 2023; Dong et al., 2023) use general LLM (ChatGPT) to get better alignment with human preferences. The Aplaca prompts (Taori et al., 2023) is used as sampling queries to different models generate responses. And these data was evaluated by LLM-as-a-Judge to obtain human preference scores (reward score) to train a new language model. Other works would like to use Supervised Fine-Tuning (SFT) model itself as evaluator, like generating better-aligned datasets for SFT including hindsight-modified prompts (Zhang et al., 2023a; Liu et al., 2023c) and principle-driven self-alignment (Sun et al., 2023c).
@@ 2.4.3. LLM-as-a-Judge for Agents
2.4.4. LLM-as-a-Judge for Reasoning/Thinking
Reasoning (Huang and Chang, 2023), defined as the cognitive process of applying logic, arguments, and evidence to draw conclusions, is central to intellectual tasks such as decision-making, problem-solving, and critical analysis. While reasoning is inherently more demanding and multifaceted than judging, it often depends on judgments to ensure logical coherence, refine intermediate steps, and achieve clarity in its outcomes. LLM-as-a-Judge, in this sense, becomes an integral tool for enhancing the reasoning capability of LLM.
The role of LLM-as-a-Judge in enhancing reasoning or thinking can be understood through two frameworks: scaling training time (Gao et al., 2023b; Trung et al., 2024) and scaling test time (Snell et al., 2024). In the training phase, LLM-as-a-Judge frequently operates within reinforcement learning paradigms, where it functions as a reward model or evaluator for data or processes. This enables the creation of high-quality reasoning datasets through mechanisms such as step-by-step verification (Lightman et al., 2023), Direct Preference Optimization(DPO) (Rafailov et al., 2024), and self-refinement (Yuan et al., 2024). Recently, several LLMs trained with reinforcement learning to exhibit advanced reasoning and thinking abilities have gained attention, such as o1222https://openai.com/index/learning-to-reason-with-llms/, DeepSeek-R1333https://api-docs.deepseek.com/news/news1120,gemini-thinking [^4] https://ai.google.dev/gemini-api/docs/thinking-mode, and QVQ [^5] https://huggingface.co/Qwen/QVQ-72B-Preview. In the test-time framework, LLM-as-a-Judge is crucial for evaluating and selecting the best reasoning paths. For example, in ”Best-of-N” generation scenarios, where multiple reasoning outputs are produced, the judge determines the most accurate and coherent response. This dual role in both training and test phases demonstrates the indispensable nature of LLM-as-a-Judge in enhancing reasoning systems.
@@ 3. Improvement Strategy
When directly utilizing LLMs to conduct evaluation tasks such as scoring, selection, pairwise comparison, or ranking, the inherent biases of LLMs like length bias, position bias, and concreteness bias(Park et al., 2024) will lead to poor evaluation results. Addressing these inherent biases and improving the overall evaluation performance of LLMs is a critical challenge for applying LLMs as evaluators. In this section, we introduce three improvement strategies aimed at enhancing the evaluation performance of LLM-as-a-judge: design strategy of evaluation prompts (in-context learning based), improvement strategy of LLMs’ evaluation capabilities (model-based), and optimization strategy of final evaluation results (post-processing based). Our categorization is based on the formal definition of LLM-as-a-judge in Section 2, focusing on enhancing the evaluation effectiveness by targeting three key phases of the process: the context $\mathcal{C}$, the abilities of LLMs themselves $\mathcal{P_{LLM}}$ and the post-processing $\leftarrow$ to obtain the final results $\mathcal{E}$
@@
Output Forms, text width=6em [Constraining outputs in structured formats: G-Eval (Liu et al., 2023a) / DHP (Wang et al., 2024d) / LLM-EVAL (Lin and Chen, 2023)] [Providing evaluations with explanations: CLAIR (Chan et al., 2023) / FLEUR (Lee et al., 2024)] ] ] [Improvement Strategy
@@
Based on Feedbacks, text width=7em [INSTRUCTSCORE (Xu et al., 2023) / JADE (Zhang et al., 2023b)] ] ] [Optimization Strategy
@@
Evaluation Results, text width=7.5em [Summarize by multiple rounds: Sottana et al. (Sottana et al., 2023) / PsychoBench (Huang et al., 2023) / Auto-J (Li et al., 2023b)] [Vote by multiple LLMs: CPAD (Liu et al., 2023d) / Bai et al. (Bai et al., 2023b)] ] [Direct Optimization of
LLMs’ Outputs, text width=7.5em [Score smoothing: FLEUR (Lee et al., 2024) / G-Eval (Liu et al., 2023a) / DHP (Wang et al., 2024d)] [Self validation: TrueTeacher (Gekhman et al., 2023)] ] ] ]